Solid lithium hydroxide is used in space vehicles to remove exhaled carbon dioxide from the living environment to form solid lithium carbonate and liquid water. What mass of gaseous carbon dioxide can be absorbed by 1.00kg of lithium hydroxide?

2LiOH(s) + CO2(g) → Li2CO3(s) + H2O(l)
1
Expert's answer
2020-05-11T14:23:57-0400

2LiOH(s) + CO2(g) → Li2CO3(s) + H2O(l)

According to the reaction, the number of moles of the reactants must be equal. As a result, the mass of carbon dioxide can be calculated from the following equation:

m(LiOH) / [2 × Mr(LiOH)] = m(CO2) / Mr(CO2)

From here:

m(CO2) = m(LiOH) × Mr(CO2) / [2 × Mr(LiOH)] = 1.00 kg × 44 g/mol / [2 × 24 g/mol] = 0.917 kg = 917 g


Answer: 917 g